Calculate [H+] for each of the following solutions, and indicate whether the solution is acidic, basic, or neutral. 1.[OH−]= 8.0×10−9 M . Express your answer using two significant figures. 2.A solution in which [OH−] is 100 times greater than [H+]. Express your answer using two significant figures. Can someone show every step to do this by,
Answer #1

1.[OH]= 8.0×10−9 M

Since,

[H+] [OH-] = 1.0 x 10-14

\Rightarrow[H+] = (1.0 x 10-14) / [OH-]

\Rightarrow[H+] = (1.0 x 10-14) / (8.0×10−9)

\Rightarrow[H+] = 1.25 x 10-6

\Rightarrow[H+] = 1.3 x 10-6

2. [OH-]= 100 [H+]

[H+][OH-] = 1.0 x 10-14

[H+] x 100 [H+] = 1.0 x 10-14

100 [H+]2 = 1.0 x 10-14

[H+]2 = 1.0 x 10-16

[H+] = \sqrt{}(1.0 x 10-16) =1.0 x 10-8 M